9x^2 +25y^2 = 64 into form of (x-h)^2/a^2 + (y-k)^2/b^2, How?

Answer by lwsshak3(6490)

9x^2 +25y^2 = 64 into form of (x-h)^2/a^2 + (y-k)^2/b^2, How?
..
9x^2 +25y^2 = 64
x^2/(64/9)+y^2/(64/25)=1
This is an ellipse with center at (0,0) and horizontal major axis
Standard form of ellipse: (x-h)^2/a^2+(y-k)^2/b^2=1
For given equation:
a^2=64/9
a=8/3
b^2=64/25
b=8/5
center: (h,k)=(0,0)
The graph below shows what this ellipse looks like
..
y=((64-9x^2)/25)^.5